whats the equation that tells how far someone will be from home if they started out 50 miles from home and they are driving a constant speed of 45 mph?
Answer by Mathtut(3670)

let x be the amount of miles driven and t= time
:
50-x=45(t) if you drive for t=1 hour(time has to be in hours) because the distance is in miles and the speed is mph. that would be 45 miles
:
50-x=45(1)-->x=5...so 5 miles from home